OpenStudy (anonymous):

a rectangle has a length of (x-3) and a width of (3x^2+4x). What is the perimeter?

OpenStudy (angela210793):

P=2(x-3)+2(3x^2+4x)=2x-6+6x^2+8x=6x^2+10x-6

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

simpler forms- 2(3x^2+5x-3) 2(3x-1)(x+3) same answer but simplified